OpenStudy (anonymous):

Find the derivative of f(x) = 3x + 8 at x = 4.

OpenStudy (anonymous):

using short cuts f'(x) = 3 therefor, f'(4) = 3. using the definition of a derivative \[f'(x) = (f(x + h) - f(x))/(h)\] \[f(x) = 3x + 8\] \[f(x+h) = 3(x + h) + 8\] \[f(x+h) = 3x + 3h + 8\] \[f(x+h)-f(x) = 3x + 3h + 8 - 3x - 8\] \[f(x+h)-f(x) = 3h\] \[f'(x) = 3h/h\] \[f'(x) = 3\]